OpenStudy (anonymous):

terminal side of an angle theta in standard position passes through the point (2,5). Find sec theta

OpenStudy (anonymous):

\[\theta=\tan^{-1} (\frac{ y }{ x })=\tan^{-1} (\frac{ 5 }{ 2 })\]

OpenStudy (anonymous):

now can you find theta ? @martaamador62

OpenStudy (anonymous):

yes thank you :)

OpenStudy (anonymous):

it's 68.20 degrees

OpenStudy (anonymous):